# Sommerfeld Enhancement for Vector Dark Matter in the Fundamental Representation of $SU(2)_{L}$

### v1

@@ Abstract
In this work, we study an extension of the Standard Model which include a new massive vector field in the fundamental representation of $SU(2)_{L}$. The neutral component of the new vector field is a natural dark matter candidate. We compute the Annihilation Cross Section with Sommerfeld Enhancement and the Photon Flux arising from Dark Matter annihilation events. We found that the model exhibits a resonant mass part around the range of $5-10$ TeV. This resonance is influenced by the choice of portal Higgs coupling parameter. We show that these prediction can be partially tested by CTA in the near future
@@ 1 Introduction
Currently, one the most urgent problems in Particle Physics revolves around the elucidation of Dark Matter’s (DM) [2] fundamental nature. It is well known that,to account for the overarching phenomenology of DM, it suffices to posit the existence of a new particle falling within a mass range spanning from a few GeV to a few TeV, characterized by an interaction strength with the standard sector on the order of the weak interaction. This phenomenon is commonly referred to as the ”WIMP miracle,” with WIMP representing ”Weakly Interacting Massive Particle [12].
An intriguing realization of the WIMP idea is to postulate the DM as the neutral component of an electroweak multiplet. This paradigm, usually called “Minimal Dark Matter”, was proposed long ago [8] and have been studied in datail for scalar field [7] and fermionic fields [10].
In the case of vector DM (VDM), an effective model including a VDM candidate originated from a vector field in the adjoint representation of $SU(2)_{L}$ was first proposed in [4] while an ultraviolet complete model was developed in [1]. On the other had, in the fundamental representation, models including a vector DM candidate coming from an electroweak doublet were studied in  [15, 18]. These vector models are consistent with experimental data, if the new fields has mass in the range of some few TeV. In this mass range, direct detection [19] becomes difficult and the discovery at colliders seems to be challenging even for future colliders [4].
In this circumstances, indirect search can be a privileged alternative to discover or rule out these models. However, it is well known that to obtain an accurate value of the DM annhihilation cross section in the non-relativistic regime, it is necessary to take into account non-perturbative effects like the so called Sommerfeld Enhancement (SE).
The SE effect was computed for the cases of scalar DM field [3, 7, 9, 11, 13] showing that effect has a resonant part, fermion DM field [5, 10] in its doublet and triplet representation of $SU(2)_{L}$ and vector DM in the adjoint representation in  [1]. In this work, we focus on computing the SE for the annihilation of VDM in the fundamental representation of $SU(2)_{L}$ and the photon flux (relevant for Indirect Detection  [6, 14, 19]) produced by pair annihilation of the VDM.
Our work is organized as follows. In section 2, we describe the main aspects of the model. Section 3 is devoted to the computation of the annihilation matrices. In section 4 we apply our results to compute the photon flux from DM annihilation at the center of the Milky Way and describe our results. Finally, in section 5 we state our conclusions.
@@ 2 The Model
We start by recalling the main characteristics of the model initially developed in reference  [15]. In this context, the Standard Model (SM) is extended by the introduction of a vector field $V_{\mu}$ which transforms as (1,2,1/2) under the standard group $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$ This new vector field can be explicitly written as:
@@
In principle, the two neutral states $V_{\mu}^{1}$ and $V_{\mu}^{2}$ (the vector and axial-vector fields) can have different masses. However, the requirement of unitarity preservation at low energies constrains the parameter space of the model in such a way that $V_{\mu}^{1}$ and $V_{\mu}^{2}$ become degenerated or nearly degenerated. In this work, we assume that $V_{\mu}^{1}$ and $V_{\mu}^{2}$ are exactly degenerated. In this case they combine into a complex neutral field $V_{\mu}^{0}$.
@@
where $B^{\mu\nu}$ represents the Abelian strength field of $U(1)_{Y}$ and $W^{\mu\nu}=W^{\mu\nu~{}a}\dfrac{\tau^{a}}{2}$ represents the non-Abelian strength field of $SU(2)_{L}$. Notice that the las two terms of Lagrangian (2) describe non-minimal interactions between the new vector field and the gauge bosons. In order to avoid a direct photon interaction with $V_{\mu}^{0}$, the otherwise arbitrary constants $\kappa_{1}$ and $\kappa_{2}$ have to be equal. In this work, however, for the sake of simplicity and minimality we assume that $\kappa_{1}=\kappa_{2}=0$. Besides, the coupling constant that regulates the interactions between the Higgs field and the new vector field can be treated a free parameters. For simplicity, we will take the assumption $\lambda_{2}=\lambda_{3}=\lambda_{4}=\lambda$). In this way, $\lambda$ will be the only constant regulating the Higgs portal.
Due to a fortunate conspiracy between the Lorentz and gauge structures of the model, the new vector field can only appear in pairs. That means that the Lagrangian is invariant by a $Z_{2}$ transformation under which the new vector is odd and all the standard fields are even. Consequently, this accidental $Z_{2}$ symmetry makes the lighter component of $V_{\mu}$ stable. This stable component happens to be the neutral one ($V_{\mu}^{0}$)
@@ 3.1 Basic Formalism
The so called “Sommerfeld enhancement” is a non-perturbative effect due to the resumation of initial states interactions. This effect may produce sensible consequences for the annihilation cross section in the non-relativistic regime  [10, 17]. In the few next paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect and the we will apply it to our model.
@@
$\left[\left(i\partial_{R^{0}}+\frac{\nabla_{R}^{2}}{4m}+\frac{\nabla_{r}^{2}}{m}\right)-V(r)+i\frac{N}{2}\Gamma\delta^{3}(r)\right]$ $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ (7)
@@
Here, $m$ represents the mass of Dark Matter, $\phi_{i}$ represents the DM field, $N$ is the number spin states for the two bodies system (for spin 1, we have $N=9$ ) and $\Gamma$ is the so called annihilation matrix. The coordinate $=(R^{0},\textbf{R})$ R ) denotes the center-of-mass coordinate, and r represents the relative coordinate of the Dark Matter multiplets, defined as:
@@
The imaginary part of the diagonal component of this Green function $\bra{0}T\phi_{b}\phi_{a}^{\dagger}\ket{0}$, is associated with the total annihilation cross section. We proceed to decompose the Green functions into individual partial wave modes
@@
$E=P^{0}-\dfrac{\textbf{P}}{4m}=\dfrac{mv^{2}}{4}$ (10)
@@
$\left[-\dfrac{1}{mr}\dfrac{d^{2}}{dr^{2}}r-\dfrac{l(l+1)}{mr^{2}}-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]G_{ba}^{(E,l)}(r,r^{\prime})=\dfrac{\delta(r-r^{\prime})}{r^{2}}\delta_{ba}$ (11)
In this analysis, we focus on the annihilation processes with orbital angular momentum $l=0$ (s-wave), which provide the leading-order contributions in the non-relativistic limit. Consequently, the total angular momentum $J$ is equal to the total spin angular momentum $S$. The leading-order potential terms exhibit spherical symmetry, resulting in the conservation of $J$ in annihilation processes. Furthermore, the conservation of total electric charge $Q$ is upheld. As a result, the Green function is expressed in the diagonal form for each eigenstate corresponding to $Q$ and $J$. To simplify the differential equation, we introduce $G_{ba}^{(E,l)}(r,r^{\prime})=\frac{g_{ba}(r,r^{\prime})}{rr^{\prime}}$.
$\left[-\dfrac{1}{mr}\dfrac{d^{2}}{dr^{2}}r-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]g_{ba}^{(E,l)}(r,r^{\prime})=\delta(r-r^{\prime})\delta_{ba}$ (12)
The cross-section is determined by the imaginary part of the solution, and we derive the leading-order solution within the perturbative expansion of $\alpha$. The leading order solution can be written as
@@
$\left[-\dfrac{1}{m}\dfrac{d}{dr^{2}}-E+V(r)\right]g_{>ab}(r)=0$ (14)
with the boundary conditions
@@
$g_{>ba}(r)\longrightarrow\ \left\{\begin{array}[]{lcc}d_{ba}(E_{a})e^{i\sqrt{mE_{b}}r}\hskip 28.45274ptshort-range~{}potential\\ \\ d_{ba}(E_{a})e^{i\sqrt{mE_{b}}r}e^{i\frac{m\alpha}{2\sqrt{mE_{b}}}\log{2\sqrt{mE_{b}}r}}\hskip 28.45274ptlarge-range~{}potential\\ \end{array}\right.$ (19)
@@
We numerically solve equation (14) to obtain the Sommerfeld factor as a function of the relative velocity for the initial state $v$. This effect becomes significant in the non-relativistic regime ($v\ll 1$), where the potential energy becomes comparable to the kinetic energy ($E\simeq\frac{mv^{2}}{4}$ ) in the Schrödinger equations. Conversely, the enhancement factors approach unity for $v\gg 1$. Consequently, we can recover the perturbative cross-section, as obtained by the Born-Oppenheimer approximation, by replacing the Sommerfeld factor with 1.
@@ 3.2 Annihilation Cross Section
$\braket{\sigma_{a}v}=c_{a}\dfrac{m^{2}}{k^{2}}\lim_{E\longrightarrow\frac{k^{2}}{m^{2}}}\left(E-\dfrac{k^{2}}{m^{2}}\right)^{2}[\mathcal{A}(k)~{}\Gamma~{}\mathcal{A}^{\dagger}(k)]_{aa}$ (25)
@@
$=d_{ba}(E)\left(-\dfrac{k}{m}\right)\left(E-\frac{k^{2}}{m}\right)^{-1}$ (28)
@@
Here, $c$ is set to $2$ for initial states consisting of identical particles. We define $E$ as approximately $E\simeq\frac{mv_{rel}^{2}}{4}$ representing the non-relativistic kinetic energy of the V-particles, and $d_{\alpha\beta}(E)~{}(\alpha,\beta=1,2,3,4)$ as the Sommerfeld enhancement factor.
@@ 3.3 Non-Relativistic Limit
$V_{\mu}^{0}(x)=\dfrac{1}{\sqrt{2m}}[v_{0}(\vec{x})\epsilon_{\mu}(x)e^{imt}+a_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-imt}]\hskip 14.22636pt$ (31)
$(V_{\mu}^{0})^{\ast}(x)=\dfrac{1}{\sqrt{2m}}[a_{0}(\vec{x})\epsilon_{\mu}(x)e^{imt}+v_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-imt}]\hskip 14.22636pt$ (32)
$V_{\mu}^{+}(x)=\dfrac{1}{\sqrt{2m}}[v_{+}(\vec{x})\epsilon_{\mu}(x)e^{imt}+v_{-}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-imt}]\hskip 14.22636pt$ (33)
[…]

### v2

@@ Abstract
In this work, we study an extension of the Standard Model which include a new massive vector field in the fundamental representation of $SU(2)_{L}$. The neutral component of the new vector field is a natural dark matter candidate. We compute the Annihilation Cross Section with Sommerfeld Enhancement and the Photon Flux arising from Dark Matter annihilation events. We found that the model exhibits a resonant mass part around the range of $2-10$ TeV. This resonance is influenced by the choice of portal Higgs coupling parameter. We show that these prediction can be partially tested by CTA in the near future.
@@ 1 Introduction
Currently, one of the most urgent problems in Particle Physics revolves around the elucidation of Dark Matter’s (DM) fundamental nature  [1]. It is well known that, to account for the overarching phenomenology of DM, it suffices to posit the existence of a new particle falling within a mass range spanning from a few GeV to a few TeV, characterized by an interaction strength with the standard sector on the order of the weak interaction. This phenomenon is commonly referred to as the ”WIMP miracle”, with WIMP representing ”Weakly Interacting Massive Particle” [2].
An intriguing realization of the WIMP idea is to postulate the DM as the neutral component of an electroweak multiplet. This paradigm, usually called “Minimal Dark Matter”, was proposed long ago [3] and have been studied in detail for scalar field [4] and fermionic fields [5].
In the case of vector DM (VDM), an effective model including a VDM candidate originated from a vector field in the adjoint representation of $SU(2)_{L}$ was first proposed in [6] while an ultraviolet complete model was developed in [7]. On the other hand, models including a vector DM candidate coming from an electroweak doublet were studied in  [8, 9]. These vector models are consistent with experimental data, if the new fields has mass in the range of some few TeV. In this mass range, direct detection [10] becomes difficult and the discovery at colliders seems to be challenging even for future colliders [6].
In this circumstances, indirect search can be a privileged alternative to discover or rule out these models. However, it is well known that to obtain an accurate value of the DM annihilation cross section in the non-relativistic regime, it is necessary to take into account non-perturbative effects like the so called Sommerfeld Enhancement (SE).
The SE effect was computed for the cases of scalar DM [4, 11, 12, 13, 14], fermion DM field [5, 15] in the doublet and triplet representation of $SU(2)_{L}$ and vector DM in the adjoint representation  [7]. In this work, we focus on computing the SE for the annihilation of VDM in the fundamental representation of $SU(2)_{L}$ and the photon flux (relevant for Indirect Detection  [10, 16, 17, 18]) produced by pair annihilation of the VDM. As we will see in the following section, a unique feature of our case is the presence of non-minimal interactions between the massive vector doublet and the standard gauge bosons.
Our work is organized as follows. In section 2, we describe the main aspects of the model. Section 3 is devoted to the computation of the electroweak, annihilation matrices and annihilation cross section with SE. In section 4 we apply our results to compute the photon flux from DM annihilation at the center of the Milky Way and describe our results. Finally, in section 5 we state our conclusions.
@@ 2 The Model
We start by recalling the main characteristics of the model initially developed in reference  [8]. In this context, the Standard Model (SM) is extended by the introduction of a vector field $V_{\mu}$ which transforms as (1, 2, 1/2) under the standard group $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$. This new vector field can be explicitly written as:
@@
In principle, the two neutral states $V_{\mu}^{1}$ and $V_{\mu}^{2}$ (the vector and axial-vector fields) can have different masses. However, unitarity constrains force the neutral components $V_{\mu}^{1}$ and $V_{\mu}^{2}$ to be fully degenerate  [8]. In this case they combine into a complex neutral field $V_{\mu}^{0}$.
@@
where $B^{\mu\nu}$ represents the Abelian strength field of $U(1)_{Y}$ and $W^{\mu\nu}=W^{\mu\nu a}\frac{\tau^{a}}{2}$ is the non-Abelian strength field of $SU(2)_{L}$. The lagrangian mass of the vector field $\tilde{m}$ will be corrected after the Higgs boson acquires its vacuum expectation value, due to the interaction terms between the Higgs field and the vector field present in the potential. We will call $m$ the physical mass of the vector boson.
Notice that the last two terms of Lagrangian (2) describe non-minimal interactions between the new vector field and the gauge bosons. The appearance of this non-minimal dimension 4 terms is a special feature of having introduced a vector doublet. Similar non-minimal interactions in the scalar and fermion cases happen to be higher dimensional. In order to avoid a direct photon interaction with $V_{\mu}^{0}$, the otherwise arbitrary constants $\kappa_{1}$ and $\kappa_{2}$ have to be equal. For the purpose of comparing, we work with two values of $\kappa=\kappa_{1}=\kappa_{2}$: $\kappa=0$ and $\kappa=1$. The last value coincides with the choice made in [8].
Another important sector of the Lagrangian is the part of the potential which describes the interaction between the Higgs field and our massive vector doublet. Of course, the coupling constant that regulates the interactions between the Higgs field and the new vector field are free parameters. However, there is no reason to think that an important hierarchy arise among these couplings. Consequently, and for the sake of simplicity, we will take the assumption that all of them are of the same order $\lambda_{2}\approx\lambda_{3}\approx\lambda_{4}\approx\lambda$. In this way, $\lambda$ will be the only constant regulating the Higgs portal.
Finally, the $V_{\mu}$ quartic self-interactions will not contribute significantly to the SE since this phenomenon is dominated by interactions mediated by particles much lighter than $V_{\mu}$, like the electroweak gauge bosons and the Higgs boson. Consequently, we can set $\alpha_{2}=\alpha_{3}=0$.
Due to a fortunate conspiracy between the Lorentz and gauge structures of the model, the new vector field can only appear in pairs in the Lagrangian. That means that the Lagrangian is invariant by a $Z_{2}$ transformation under which the new vector is odd and all the standard fields are even. Consequently, this accidental $Z_{2}$ symmetry makes the lighter component of $V_{\mu}$ stable. This stable component happens to be the neutral one ($V_{\mu}^{0}$)
@@ 3.1 Basic Formalism
The so called “Sommerfeld enhancement” is a non-perturbative effect due to the resumation of initial states interactions. This effect may produce sensible consequences for the annihilation cross section in the non-relativistic regime [5, 19]. In the next few paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect and the we will apply it to our model.
@@
$\left[\left(i\partial_{R^{0}}+\frac{\nabla_{R}^{2}}{4M_{V}}+\frac{\nabla_{r}^{2}}{M_{V}}\right)-V(r)+i\frac{N}{2}\Gamma\delta^{3}(r)\right]$ $\bra{0}T\phi_{b}(R^{\prime},\textbf{r}^{\prime})\phi_{a}^{\dagger}(R,\textbf{r})\ket{0}$ (3)
@@
Here, $m$ represents the mass of Dark Matter, $\phi_{i}$ represents the DM field, $N$ is the number spin states for the two bodies system (for spin 1, we have $N=9$), $V(r)$ is electroweak matrices and $\Gamma$ is the so called annihilation matrix. The coordinate $=(R^{0},\textbf{R})$ R ) denotes the center-of-mass coordinate, and r represents the relative coordinate of the Dark Matter multiplets, defined as:
@@
The imaginary part of the diagonal component of this Green function $\bra{0}T\phi_{b}\phi_{a}^{\dagger}\ket{0}$, is associated with the total annihilation cross section. We proceed to decompose the Green functions into individual partial wave modes.
@@
$E=P^{0}-\dfrac{\textbf{P}^{2}}{4M_{V}}=\dfrac{M_{V}v^{2}}{4}$ (6)
@@
$\left[-\dfrac{1}{M_{V}r}\dfrac{d^{2}}{dr^{2}}r-\dfrac{l(l+1)}{M_{V}r^{2}}-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]G_{ba}^{(E,l)}(r,r^{\prime})=\dfrac{\delta(r-r^{\prime})}{r^{2}}\delta_{ba}$ (7)
In this analysis, we focus on the annihilation processes with orbital angular momentum $l=0$ (s-wave), which provide the leading-order contributions in the non-relativistic limit. Indeed, we verified that the $l=1$ contribution is at least one order of magnitude smaller than the leading s-wave. Consequently, in our calculations the total angular momentum $J$ is equal to the total spin angular momentum $S$. The leading-order potential terms exhibit spherical symmetry, resulting in the conservation of $J$ in annihilation processes. Furthermore, the conservation of total electric charge $Q$ is upheld. As a result, the Green function is expressed in the diagonal form for each eigenstate corresponding to $Q$ and $J$. To simplify the differential equation, we introduce $G_{ba}^{(E,l)}(r,r^{\prime})=\frac{g_{ba}(r,r^{\prime})}{rr^{\prime}}$.
$\left[-\dfrac{1}{M_{V}r}\dfrac{d^{2}}{dr^{2}}r-E+V(r)-i\Gamma\dfrac{N\delta(r)}{8\pi r^{2}}\right]g_{ba}^{(E,l=0)}(r,r^{\prime})=\delta(r-r^{\prime})\delta_{ba}$ (8)
The cross-section is determined by the imaginary part of the solution, and we derive the leading-order solution within the perturbative expansion of $\alpha_{2}$. The leading order solution can be written as
@@
$\left[-\dfrac{1}{M_{V}}\dfrac{d}{dr^{2}}-E+V(r)\right]g_{>ab}(r)=0$ (10)
Where $V(r)=V_{g-V}(r)+V_{h-V}(r)$ represents the electroweak potential, composed of $V_{g-V}(r)$ (interaction gauge bosons-VDM) and $V_{h-V}(r)$ (interaction Higgs boson-VDM) terms, $E$ denotes the energy of dark matter particles, $m$ stands for the mass of dark matter, and $g_{>ab}(r)$ denotes the wave function for $(r\gg 1)$, with the corresponding boundary conditions.
@@
$g_{>ba}(r)\longrightarrow\ \left\{\begin{array}[]{lcc}d_{ba}(E_{a})e^{i\sqrt{M_{V}E_{b}}r}\hskip 28.45274ptshort-range~{}potential\\ \\ d_{ba}(E_{a})e^{i\sqrt{M_{V}E_{b}}r}e^{i\frac{M_{V}\alpha}{2\sqrt{M_{V}E_{b}}}\log{2\sqrt{M_{V}E_{b}}r}}\hskip 28.45274ptlarge-range~{}potential\\ \end{array}\right.$ (15)
@@
We numerically solve equation (10) to obtain the Sommerfeld factor as a function of the relative velocity for the initial state $v$. This effect becomes significant in the non-relativistic regime ($v\ll 1$), where the potential energy becomes comparable to the kinetic energy ($E\simeq\frac{M_{V}v^{2}}{4}$ ) in the Schrödinger equations. Conversely, the enhancement factors approach unity for ($v\approx 1$). Consequently, we can recover the perturbative cross-section, as obtained by the Born-Oppenheimer approximation, by replacing the Sommerfeld factor with $1$.
@@ 3.2 Annihilation Cross Section
$\braket{\sigma_{a}v}=c_{a}\dfrac{M_{V}^{2}}{k^{2}}\lim_{E\longrightarrow\frac{k^{2}}{M_{V}^{2}}}\left(E-\dfrac{k^{2}}{m^{2}}\right)^{2}[\mathcal{A}(k)~{}\Gamma~{}\mathcal{A}^{\dagger}(k)]_{aa}$ (21)
@@
$=d_{ba}(E)\left(-\dfrac{k}{M_{V}}\right)\left(E-\frac{k^{2}}{M_{V}}\right)^{-1}$ (24)
@@
Here, the index $3$ must be fixed; therefore, the Sommerfeld enhancement factor implies that we are considering the process $(V_{0}V_{0}^{\ast}\longrightarrow XX^{\prime})$, with $c$ is set to $2$ for initial states consisting of identical particles. We define $E$ as approximately $(E\simeq\frac{M_{V}v_{rel}^{2}}{4})$ representing the non-relativistic kinetic energy of the V-particles, $(\Gamma_{XX^{\prime}}^{J,J_{z}})_{\alpha\beta}$ is the Annihilation matrices, with $X$ is any standard model particle, and $d_{\alpha\beta}(E)$ with $(\alpha,\beta=1,2,3,4)$ as the Sommerfeld enhancement factor.
@@ 3.3 Non-Relativistic Limit
$V_{\mu}^{0}(x)=\dfrac{1}{\sqrt{2M_{V}}}[v_{0}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}+a_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}]\hskip 14.22636pt$ (27)
$(V_{\mu}^{0})^{\ast}(x)=\dfrac{1}{\sqrt{2M_{V}}}[a_{0}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}+v_{0}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}]\hskip 14.22636pt$ (28)
$V_{\mu}^{+}(x)=\dfrac{1}{\sqrt{2M_{V}}}[v_{+}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}+v_{-}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}]\hskip 14.22636pt$ (29)
$V_{\mu}^{-}(x)=\dfrac{1}{\sqrt{2M_{V}}}[v_{+}^{\dagger}(\vec{x})\epsilon_{\mu}^{\ast}(x)e^{-iM_{V}t}+v_{-}(\vec{x})\epsilon_{\mu}(x)e^{iM_{V}t}]\hskip 14.22636pt$ (30)